40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 429     Entry time: Sun Apr 20 18:23:27 2008
Author: rana 
Type: Summary 
Category: LSC 
Subject: locking attempts 
I noticed that the adaptive FF for the MC had stopped doing anything; this turned out
to be that the MC lost lock and the mcdown script turned off the FF path to MC1.

Although there's no elog, it looks like there was ~60 attempts at locking the IFO
between 12:38 and 4:27 on Saturday afternoon. I'm attaching here a plot showing
lock attempt durations and a histogram of lock times.
Attachment 1: quix.png  16 kB  | Hide | Hide all
quix.png
ELOG V3.1.3-